OpenStudy (anonymous):

verify the identity: 1) tanx^2/secx=sinxtanx 2) cotx^2/cscx=(1-sinx^2)/sinx

OpenStudy (lalaly):

1)\[\huge{\frac{(\frac{sinx}{cosx})^2}{\frac{1}{cosx}}}\]\[=\frac{\sin^2x}{cosx}\]\[\huge{=\frac{sinx}{cosx}\times sinx}=tanxsinx\]

OpenStudy (lalaly):

same thing for 2 \[\huge{\frac{\cot^2x}{cosecx}}\]\[\huge{=\frac{(\frac{cosx}{sinx})^2}{\frac{1}{sinx}}}\]\[\huge{=\frac{\cos^2x}{sinx}}=\frac{(1-\sin^2x)}{sinx}\]

OpenStudy (anonymous):

thank you for clearing this up for me. i forgot that the exponent cancelled out with one of the denominators.

OpenStudy (lalaly):

:)